• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name ZAFAR AHMED

    Log entry time 07:52:32 on April 2,2010

    Entry number 315721

    Followups:

    keyword=April 2nd Owl Shift Summary

    12:00---Shift starts, Dustin, Paul and Kent are working on detector alignment. Beam is off, target is empty raster is 2x1 MCC.

    04:00---John moved the detectors 2cm away from the target in X.

    04:20---John moved the detectors 1/2cm towards beam beam line in both arms.

    04:38---John: Movers moved again. The left arm moved the detector another 5mm towards the beam line, so the current position is (-20mm, -10mm). The right arm moved the detector 7mm away from the beam line. The current position is therefore (-20mm, -2mm)

    05:10---Right detector moved -4mm in y (away from beamline). Current position is (-20mm, -6mm).

    05:39---Moved to empty target.

    07:31---Asked for escorted access.